### v1

# Causal Inference with Large Language Model: A Survey

Abstract
Causal inference has been a pivotal challenge across diverse domains such as medicine and economics, demanding a complicated integration of human knowledge, mathematical reasoning, and data mining capabilities. Recent advancements in natural language processing (NLP), particularly with the advent of large language models (LLMs), have introduced promising opportunities for traditional causal inference tasks. This paper reviews recent progress in applying LLMs to causal inference, encompassing various tasks spanning different levels of causation. We summarize the main causal problems and approaches, and present a comparison of their evaluation results in different causal scenarios. Furthermore, we discuss key findings and outline directions for future research, underscoring the potential implications of integrating LLMs in advancing causal inference methodologies.
Causal Inference with Large Language Model: A Survey
Jing Ma Case Western Reserve University, Cleveland, OH 44106 jing.ma5@case.edu
1 Introduction
1.1 NLP, LLM, and Causality
Causal inference is an important area to uncover and leverage the causal relationships behind observations, enabling a deep understanding of the underlying mechanism and potential interventions in real-world data systems. Different from most classical statistical studies, causal inference presents unique challenges due to its focus on "causation instead of correlation", which intricates a complicated integration of human knowledge (e.g., domain expertise and common sense), mathematics, and data mining. Due to the inherent proximity to the human cognitive process, causal inference has become pivotal in many high-stakes domains such as healthcare Glass et al. (2013), finance Atanasov and Black (2016), and science Imbens and Rubin (2015).
Traditional causal inference frameworks, such as structural causal model (SCM) Pearl (2009) and potential outcome framework Imbens and Rubin (2015) have systematically defined causal concepts, quantities, and measures, followed up with multiple data-driven methods to discover the underlying causal relationships Spirtes and Zhang (2016); Nogueira et al. (2022); Vowels et al. (2022) and estimate the significance of causal effects Winship and Morgan (1999); Yao et al. (2021). Despite their success, there is still a large gap between existing causal methods and human’s judgment Kıcıman et al. (2023); Zečević et al. (2023); Jin et al. (2023a), covering different aspects such as the lack of domain knowledge, logic inference, and cultural background. Besides, most traditional causal inference approaches only focus on tabular data, lacking the ability to discover and utilize the causality inside natural language. However, the motivation for causal inference in natural language has persisted over an extended period, offering a multitude of potential applications. For example, clinical text data in electronic health records (EHR) contains a large amount of underlying causal knowledge that can be utilized for healthcare-related research. In general, causal inference in natural language processing (NLP) is a promising research path with strong motivation, which offers a spectrum of challenges and benefits concurrently. Recently, the burgeoning field of large language models (LLMs) has shed light on its potential to improve traditional causal inference, offering fresh perspectives to bridge the gap between human cognition and causal inference methodologies Feder et al. (2022).
Figure 1: Representative causal tasks, their positions in the causal ladder, and examples of prompts. PCD = pairwise causal discovery; CA=causal attribution; ATE=average treatment effect; CDE=controlled direct effect; BAJ=backdoor adjustment; CE=causal explanation; CR=counterfactual reasoning; NDE=natural direct effect.
1.2 Challenges of Causal Inference in NLP
Although LLMs have shown eye-catching success in various tasks, causal inference still presents many distinctive challenges for LLM capabilities. Different from regular data types, the nature of natural language brings difficulties in causal processing and analysis. As aforementioned, text data is often unstructured, high-dimensional, and large-scale, in which context traditional causal methods are not applicable. Besides, causal relationships inside the text are often obscure and sparse. The complicated semantic meaning and ambiguity hidden in text data require sophisticated language modeling technologies to discover clear causal relationships, and also entail hurdles for other causal tasks such as causal intervention and counterfactual reasoning. These challenges demand new perspectives, assumptions, and technologies to address them effectively, offering revolutionary opportunities for current causal inference studies.
1.3 Opportunites that LLMs Bring to Causal Inference
Despite the challenges, natural language has significant potential to yield advantages in causal inference. As LLMs have become increasingly sophisticated with diverse applications in recent years, the feasibility of understanding and unraveling causal relationships within linguistic data has been substantially improved. In general, LLM can bring many benefits to causal inference, including but not limited to the following main aspects:
Domain knowledge. Typical statistical methods for causal inference often only focus on the numerical values of variables, while in many scenarios, domain knowledge plays an important role in causality-related tasks as it provides us with additional information to discover the true causal relationships and make meaningful interventions. For example, in many scientific domains such as medicine, incorporating the domain knowledge can draw conclusions that cannot be obtained solely through pure statistical methods, and expedite the development of relevant fields. However, collecting domain knowledge from human experts often demands considerable effort. Fortunately, the recent developments in NLP and LLM can extract domain knowledge from large-scale text information and thereby facilitate causal inference.
Common sense. Similar to domain knowledge, language models can serve as an effective tool to learn and utilize humans’ general common sense to promote causal inference. As discussed in Kıcıman et al. (2023), a variety of common sense in different scenarios affects humans’ recognition of causal relationships. For example, logical reasoning is essential for causal inference in law cases. Besides, abnormal events are often more likely to be recognized as causes for an outcome of interest in common sense.
Sematical concept. Compared with regular data types, natural language contains nuances, variations, and the richness of human expression, requiring advanced techniques for semantic analysis. Therefore, grasping clear causal concepts and relationships from text data is much more challenging than other data types. Recent progress in NLP and LLM technologies, especially their ability in semantic modeling pave the way for in-depth causal studies in the next step.
Explainable causal inference. LLMs have the potential to offer sophisticated natural language-based tools that facilitate a more engaging and intuitive human understanding of causal inference. By leveraging these tools, users can gain clearer insights into the causal reasoning process and the resulting findings, making intricate concepts more accessible. This approach not only simplifies the interpretation of causal relationships but also enhances the ability of individuals to interact meaningfully with the causal inference results, thereby bridging the gap between complex analytical processes and user comprehension.
1.4 Contribution and Uniqueness
Contribution. This survey systematically reviews existing studies of using LLMs for causal inference. The main contribution of our survey can be summarized as:
  • We provide a well-structured categorization for existing studies in this area, organizing them into distinct groups based on their tasks (Section 2) and technologies (Section 3).
  • We show a detailed evaluation comparison of different existing LLMs (Section 4), and we elucidate the main observations, connections, and insights based on the results.
  • We present a comprehensive summarization of benchmark datasets in the field, covering many important aspects for further study, as illustrated in Table 1.
  • We discuss the limitations of current studies and future directions in Section 5. These discussions highlight gaps and further opportunities that have not been fully explored in existing literature, offering a novel perspective on potential advancements in the area.
Differences from existing surveys. There have been several related surveys of LLM applications in causal inference Liu et al. (2024b); Kıcıman et al. (2023). We summarize the differences between our survey and existing ones as follows: (1) Main scope. Our paper offers an extensive and comprehensive exploration of LLMs in causal inference, i.e., “LLMs for causality", different from some of existing surveys Liu et al. (2024b) that primarily focus on “causality for LLMs". (2) Structure and content. The structure of our survey is significantly different from previous surveys, including a well-organized presentation dedicated to tasks, methods, datasets, and evaluation, thereby offering a clearer and more thorough examination. (3) Up-to-date. Our survey incorporates cutting-edge research developments, providing an up-to-date snapshot of the latest progress and trends.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-black, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=4pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1 text width=7.0em, font=, inner xsep=4pt, , where level=2 text width=11.0em, font=, inner xsep=5pt, , where level=3 text width=12em, font=, inner xsep=4pt, , where level=4 text width=15em, font=, inner xsep=4pt, , [ LLMs for Causal Inference, ver [ Causal Discovery
4.2) [ Pairwise Causal Discovery [ GPT-4 (Jin et al.; Cai et al.; Liu et al.; Gao et al.; Zečević et al.;
Kıcıman et al.; Chen et al.), GPT-4 turbo (Cai et al.),
GPT-3.5 (Liu et al.; Jin et al.; Cai et al.; Gao et al.; Kıcıman et al.; Chen et al.),
GPT-3 (Long et al.; Gao et al.; Zečević et al.; Kıcıman et al.),
Ada, Babbage, Curie, Davinci, Text-ada-001, Text-babbage-001, Text-curie-001,
Text-davinci-001, Text-davinci-002, Text-davinci-003 (Chen et al.),
LLaMa2-13B (Cai et al.; Chen et al.), Claude 2 (Cai et al.),
Alpaca-7B (Jin et al.), FLAN-T5 (Liu et al.),
BERT-Base, RoBERTa-Base (Gao et al.),
Luminous (Zečević et al.), OPT (Zečević et al.; Liu et al.),
Mistral-7B (Cai et al.), Mistral-7B+Fine-tuning (Cai et al.) , leaf, text width=34em ] ] [ Full Graph Discovery
] ] [ Causal Effect
Estimation (§4.3) [ Causal Effect in Data [ Alpaca (Jin et al.), LLaMa (Jin et al.; Chen et al.), GPT-4 (Jin et al.; Chen et al.),
Ada, Babbage, Curie, Davinci, Text-ada-001, Text-babbage-001, Text-curie-001,
Text-davinci-001, Text-davinci-002, Text-davinci-003 (Chen et al.),
GPT-3.5, GPT-3, GPT-3 Instruct, GPT-4+CausalCoT (Jin et al.),
GPT-3.5 turbo+RAG (Pawlowski et al.) , leaf, text width=34em ] ] [ Causal Effect in Model [ GPT-2, GPT-Neo, GPT-3 Instruct, LLaMA, Alpaca (Stolfo et al.) , leaf, text width=34em ] ] ] [ Other Causal
Tasks (§4.4) [ Causal Attribution [ GPT-4 turbo (Cai et al.), GPT-4 (Cai et al.; Chen et al.),
GPT-3.5 (Liu et al.; Cai et al.; Chen et al.),
Ada, Babbage, Curie, Davinci, Text-ada-001, Text-babbage-001, Text-curie-001,
Text-davinci-001, Text-davinci-002, Text-davinci-003 (Chen et al.),
OPT-1.3B, FLAN-T5 (Liu et al.), LLaMa2-13B, Claude 2 (Cai et al.) , leaf, text width=34em ] ] [ Counterfactual Reasoning [ GPT-4 (Kıcıman et al.; Jin et al.; Chen et al.), GPT-3.5 (Kıcıman et al.; Jin et al.,
Chen et al.), Ada, Babbage, Curie, Davinci, Text-ada-001, Text-babbage-001,
Text-curie-001, Text-davinci-001, Text-davinci-002, Text-davinci-003 (Chen et al.),
GPT-3 (Kıcıman et al.; Jin et al.), GPT-3 Instruct (Jin et al.), T0pp (Kıcıman et al.),
Alpaca (Jin et al.), LLaMa (Jin et al.; Chen et al.), GPT-4+CausalCoT (Jin et al.) , leaf, text width=34em ] ] [ Causal Explanation [ GPT-4, GPT-3.5 (Gao et al.; Chen et al.), GPT-3 (Gao et al.; Ho et al.),
GPT-2 (Gao et al.; Ho et al.), Ada, Babbage, Curie, Davinci, Text-ada-001,
Text-babbage-001, Text-curie-001, Text-davinci-001, Text-davinci-002,
Text-davinci-003 (Chen et al.), LLaMA 7B (Gao et al.; Chen et al.),
FLAN-T5 11B (Gao et al.) , leaf, text width=34em] ] ] ]
Figure 2: The major causal tasks and LLMs evaluated for these tasks. Noticeably, the citations in the figure correspond to the work of evaluations, rather than the original work of these models themselves.
2 Preliminaries
2.1 Causality
Structural causal model. Structural causal model (SCM) Pearl (2009) is a widely used model to describe the causal relationships inside a system. An SCM is defined with a triple $(U,V,F)$: $U$ is a set of exogenous variables, whose causes are out of the system; $V$ is a set of endogenous variables, which are determined by variables in $U\cup V$; $F=\{f_{1}(\cdot),f_{2}(\cdot),...,f_{|V|}(\cdot)\}$ is a set of functions (a.k.a. structural equations). For each $V_{i}\in V$, $V_{i}=f_{i}(pa_{i},U_{i})$, where “$pa_{i}\subseteq V\setminus{V_{i}}$” and “$U_{i}\subseteq U$" are variables that directly cause $V_{i}$. Each SCM is associated with a causal graph, which is a directed acyclic graph (DAG). In the causal graph, each node stands for a variable, and each arrow represents a causal relationship.
Ladder of causation. The ladder of causation Pearl and Mackenzie (2018); Bareinboim et al. (2022) defines three rungs (Rung 1: Association; Rung 2: Intervention; Rung 3: Counterfactuals) to describe different levels of causation. Each higher rung indicates a more advanced level of causality. The first rung "Association" involves statistical dependencies, related to questions such as "What is the correlation between taking a medicine and a disease?". The second rung "Intervention" moves further to allow interventions on variables. Questions related to this rung are like "If I take a certain medicine, will my disease be cured?". The top rung "Counterfactuals" relates to imagination or retrospection queries like "What if I had acted differently?", "Why?". Answering such questions requires knowledge related to the corresponding SCM. Counterfactual ranks the highest because it subsumes the first two rungs. A model that can handle counterfactual queries can also handle associational and interventional queries.
2.2 Causal Tasks and Related Rungs in Ladder of Causation
Causal inference involves various tasks. Figure 1 shows an overview of causal inference tasks and their positions in the ladder of causation. We also show several examples of prompts corresponding to each rung. We list several main causal tasks which are most widely studied as follows:
Causal discovery. Causal discovery aims to infer causal relationships from data. It includes discovering the causal graph and the structural equations associated with these causal relationships. Although causal discovery is not explicitly covered in the ladder of causation, it is often considered as "Rung 0" as it serves as a fundamental component in causal inference. Typical causal discovery questions include pairwise causal discovery (PCD) that only focuses on a pair of variables, and full graph discovery involving variables in the whole data system.
Causal effect estimation. Causal effect estimation (a.k.a. treatment effect estimation) targets on quantifying the strength of the causal influence of a particular intervention or treatment on an outcome of interest. In different scenarios, researchers may focus on the causal effect of different granularities, such as individual treatment effect (ITE, i.e., treatment effect on a specific individual), conditional average treatment effect (CATE, i.e., average treatment effect on a certain subgroup of population), average treatment effect on the treated group (ATT), and average treatment effect (ATE, i.e., average treatment effect on the entire population). Besides, people are also interested in the direct or indirect causal effects in certain scenarios, such as natural direct effect (NDE), controlled direct effect (CDE), and natural indirect effect (NIE). Another task related to causal effect estimation is backdoor adjustment (BAJ), which aims to block all backdoor paths Pearl (2009) from the treatment to the outcome to exclude non-causal associations. Causal effect estimation tasks often span over Rung 2 and Rung 3 in the ladder of causation.
Other tasks. There are many other tasks in causal inference. Among them, causal attribution (CA) refers to the process of attributing a certain outcome to certain events. Counterfactual reasoning (CR) investigates what might have happened if certain events or conditions had been different from what actually occurred. It explores hypothetical scenarios by considering alternative outcomes based on changes in “what if" circumstances. Causal explanation (CE) aims to generate explanations for an event, a prediction, or any causal reasoning process. This task often needs to answer causal questions in a specified human-understandable form or plain language. It is often in Rung 2 or Rung 3, depending on the specific context. It is worth mentioning that, in many cases, different causal tasks may exhibit natural overlap in their scope. For instance, causal attribution and explanation commonly intersect with causal discovery and causal effect estimation. However, each task maintains a distinct focus and emphasis.
Dataset Task Size (Unit) Domain Real # of sources Citations
CEPairs (2016) CD 108 (P) Mixed R 37 (2016; 2023; 2023; 2023)
Sachs (2024) CD 20 (R) Biology R 1 (2023; 2024; 2024)
Corr2Cause (2023b) CD 200K (S) Mixed S 1 (2023b)
CLADDER (2023a) Eff, CR, CE 10K (S) Mixed S 1 (2023a; 2023b)
BN Repo (2022) CD 4$\sim$84 (R) Mixed R 8 (2023)
COPA (2011) CD 1,000 (Q) Dailylife R 1 (2023; 2011)
E-CARE (2022) CD, CE 21K (Q) Mixed R 1 (2023; 2022)
Asia (1988) CD 8 (R) Health R 1 (1988; 2024; 2024)
CausalNet (2016) CD 62M (R) Mixed S 1 (2016; 2022)
CausalBank (2021) CD 314 M (P) Mixed S 1 (2021; 2022)
WIKIWHY (2022) CD,CE 9K (Q) Mixed R 1 (2022)
Neuro Pain (2019) CD 770 (R) Health S 1 (2019; 2023; 2023)
Arctic Ice (2021) CD 48 (R) Climate R 1 (2021; 2023)
CRASS (2022) CR 275 (Q) Mixed R 1 (2022)
CausalQA (2022) CD, CE 1.1M (Q) Mixed R 10 (2024; 2022; 2024)
CALM-Bench (2023) CD, CA 281K (Q) Mixed R 6 (2023)
CausalBench (2024b) Corr, CD 4$\sim$195 (R) Mixed R 15 (2024b)
CaLM (2024a) Rung 1$\sim$3 126K (S) Mixed S 20 (2024a)
Table 1: Datasets for LLM-related causal inference, including publication year, applicable tasks (CD=causal discovery; Eff=causal effect estimation; CR=counterfactual reasoning; CE=causal explanation), dataset size (as different datasets are not in a consistent form, we show the size w.r.t. different units, where P=causal pairs; R=causal relations; S=samples; Q=questions), domain, generation process (R: real-world; S: synthetic), number of data sources, and citations.
3 Methodologies
Recently, there have emerged many efforts (Kıcıman et al., 2023; Chen et al., 2024a; Gao et al., 2023) to leverage LLMs for causal tasks. Different from traditional causal inference approaches which are either data-driven or based on expert knowledge, the nature of LLM training and adoption introduces novel methodologies in causal inference, offering new perspectives and insights for discovering and utilizing causal knowledge in future research and applications. A list of LLMs developed or evaluated in different causal tasks is shown in Figure 2. We summarize the current methodologies of LLMs for causal tasks into the following categories:
Prompting. Most existing works Chen et al. (2024a); Kıcıman et al. (2023); Long et al. (2023); Jin et al. (2023a) of causal reasoning with LLMs focus on prompting, as it is the most straightforward approach. This line of work includes both regular prompting strategies (such as basic prompt, In-Context Learning (ICL) Brown et al. (2020), and Chain-of-Thought (CoT) Wei et al. (2022)) and causality-specific strategies. For regular prompting, most studies directly use a basic prompt (i.e., directly describe the question without any example or instruction). There are also other efforts to devise more advanced prompting strategies. Among them, CaLM Chen et al. (2024a) has tested 9 prompting strategies including basic prompt, adversarial prompt Wallace et al. (2019); Perez and Ribeiro (2022), ICL, 0-shot CoT (e.g., “let’s think step by step” without any examples) Kojima et al. (2022), manual CoT (i.e., guide models with manually designed examples), and explicit function (EF) (i.e., using encouraging language in prompts) Chen et al. (2024a). Other works Kıcıman et al. (2023); Long et al. (2023); Gao et al. (2023); Ban et al. (2023) also design different prompt templates. These works show substantial improvement potential of prompt engineering in causal reasoning tasks. For example, results in Kıcıman et al. (2023); Chen et al. (2024a); Long et al. (2023) show adding simple sentences like "you are a helpful causal assistant" or "you are an expert in [DOMAIN NAME]" can impressively improve the causal inference performance for many models. Apart from these regular methods, other studies propose causality-specific prompting strategies. For example, CausalCoT Jin et al. (2023a) is a multi-step prompting strategy that combines CoT prompting and the causal inference engine Pearl and Mackenzie (2018).
Fine-tuning. Fine-tuning, as a widely recognized technique in general LLMs, is now also starting to gain attention for its application in causal tasks. Cai et al. Cai et al. (2023) propose a fine-tuned LLM for the pairwise causal discovery task (PCD, introduced in Section 4.2). This method generates a fine-tuning dataset with a Linear, Non-Gaussian, Acyclic Model Shimizu et al. (2006), uses Mistral-7B-v0.2 Jiang et al. (2023) as LLM backbone, and runs instruction finetuning with LoRA Hu et al. (2021). The results achieve significant improvement compared with the backbone without fine-tuning.
Combining LLMs with traditional causal methods. Another line of works combine LLMs with traditional causal methods. Considering causal inference often heavily relies on numerical reasoning, an exploration in Ban et al. (2023) leverages LLMs and data-driven causal algorithms such as MINOBSx Li and Beek (2018) and CaMML O’Donnell et al. (2006). This method outperforms both original LLMs and data-driven methods, indicating a promising future for combining the language understanding capability of LLMs and the numerical reasoning skills of data-driven methods in complicated causal tasks. Jiralerspong et al. Jiralerspong et al. (2024) combine LLM with a breadth-first search (BFS) approach for full causal graph discovery. It considers each causal relation query as a node expansion process, and gradually constructs the causal graph by traversing it using BFS. This method significantly reduces the time complexity from $O(n^{2})$ to $O(n)$, where $n$ is the number of variables. While it does not require access to observational data, their experiments show that the performance can be further enhanced with observational statistics.
Knowledge augmentation. LLMs with augmented knowledge can often better execute tasks for which they are not well-suited to perform by themselves. This idea is tempting for causal tasks as they often require corresponding causal knowledge. Pawlowski et al. (2023) propose two types of knowledge augmentation to answer causal questions. Their augmentations include a context augmentation that provides a causal graph and the ITEs in the prompt, and a tool augmentation with API access to an expert system that performs causal reasoning tasks. These augmented LLMs have promising results in answering causal queries. The tool augmentation performs more robustly over problems with different sizes, likely because LLM does not need to reason through the graph itself, but can instead utilize the API for causal reasoning. LACR Zhang et al. (2024) applies retrieval augmented generation (RAG) to enhance the knowledge base of LLM for causal discovery tasks, where the knowledge resources are from a large scientific corpus containing hidden insights about associational/causal relationships. Similarly, Samarajeewa et al. (2024) use causal graphs as external sources for causal reasoning. Noticeably, knowledge augmentation is particularly useful for causal tasks in professional domains. For example, RC2R Yu et al. (2024) combines LLMs with expert knowledge in financial knowledge graphs to analyze the causal mechanisms of financial risk contagion. CausalKGPT Zhou et al. (2024a) is a causal knowledge graph augmented LLM, designed to handle causal reasoning for quality defects in aerospace product manufacturing.
4 Evaluations of LLMs in Causal Tasks
4.1 Overview
In this section, we summarize recent evaluation results of LLMs in causal tasks. We mainly focus on causal discovery and causal effect estimation, and also introduce several representative tasks spanning Rung 1$\sim$3. A collection of datasets used in LLM-related causal tasks is shown in Table 1. In Table 2 and Table 3, we compare the performance of different LLMs in different tasks (including causal discovery and other tasks spanning different rungs in causal ladder) on multiple datasets. The mentioned LLMs include ada, babbage, curie, davinci Brown et al. (2020), text-ada-001, text-babbage-001, text-curie-001, text-davinci-001, text-davinci-002, text-davinci-003 Ouyang et al. (2022), Llama 2 (7B, 13B, 70B) Touvron et al. (2023), and OpenAI’s GPT series Achiam et al. (2023); OpenAI (2022).
4.2 LLM for Causal Discovery
Model CEPairs E-CARE COPA CALM-CA Neuro Pain
Binary Choice Binary Choice Binary Binary Choice
ada 0.50 0.48 0.49 0.49 0.49 0.57 0.40
text-ada-001 0.49 0.49 0.33 0.50 0.35 0.48 0.50
Llama2 (7B) - 0.53 0.50 0.41 0.35 0.32 -
Llama2 (13B) - 0.52 0.50 0.44 0.36 0.42 -
Llama2 (70B) - 0.52 0.44 0.50 0.45 0.49 -
babbage 0.51 0.49 0.36 0.49 0.40 0.58 0.50
text-babbage-001 0.50 0.50 0.50 0.49 0.50 0.56 0.51
curie 0.51 0.50 0.50 0.50 0.50 0.58 0.50
text-curie-001 0.50 0.50 0.50 0.51 0.50 0.58 0.50
davinci 0.48 0.50 0.49 0.50 0.51 0.58 0.38
text-davinci-001 0.50 0.50 0.50 0.50 0.50 0.52 0.50
text-davinci-002 0.79 0.66 0.64 0.80 0.67 0.69 0.52
text-davinci-003 0.82 0.77 0.66 0.90 0.77 0.80 0.55
GPT-3.5-Turbo 0.81 0.80 0.66 0.92 0.66 0.72 0.71
GPT-4 - 0.74 0.68 0.90 0.80 0.93 0.78
GPT-4 (0-shot ICL) - 0.83 0.71 0.97 0.78 0.90 -
GPT-4 (1-shot ICL) - 0.81 0.70 0.93 0.76 0.90 -
GPT-4 (3-shot ICL) - 0.71 0.70 0.80 0.81 0.91 -
GPT-4 (0-shot CoT) - 0.77 0.68 0.91 0.79 0.92 -
GPT-4 (Manual CoT) - 0.79 0.73 0.97 0.82 0.95 -
GPT-4 (EF) - 0.83 0.71 0.98 0.80 0.92 0.84
Table 2: Performance (accuracy) of different models in causal discovery tasks on different datasets, including CausalEffectPairs (CEpairs for short), E-CARE, COPA, CALM-CA, and Neuro Pain. In the columns in white (CausalEffectPairs, E-CARE, COPA), the models are evaluated for the pairwise causal discovery task; In the column in gray, the models are evaluated for the causal attribution task; in the column in cyan, the models are evaluated for the full graph discovery task. In the upper part, we show results with basic prompt; while in the lower part, we show results of GPT-4 with different prompting strategies. We also present results under prompts in the form of binary "yes/no" questions and multi-choice questions. The results are collected from Kıcıman et al. (2023) and Chen et al. (2024a). Note that the experimental settings such as prompt templates may be different.
Model CLADDER CaLM CLADDER CaLM CRASS E-CARE
Corr ATE CDE BAJ CR NDE CR CE
ada 0.26 0.02 0.03 0.13 0.30 0.05 0.26 0.22
text-ada-001 0.25 0.01 0.01 0.29 0.28 0.01 0.24 0.33
Llama2 (7B) 0.50 0.03 0.02 0.18 0.51 0.03 0.11 0.42
Llama2 (13B) 0.50 0.01 0.01 0.19 0.52 0.02 0.20 0.39
Llama2 (70B) 0.51 0.09 0.09 0.13 0.52 0.13 0.17 0.42
babbage 0.39 0.03 0.04 0.15 0.31 0.06 0.26 0.24
text-babbage-001 0.35 0.04 0.04 0.34 0.32 0.07 0.28 0.37
curie 0.50 0.01 0.04 0.23 0.49 0.01 0.22 0.30
text-curie-001 0.50 0.00 0.09 0.40 0.49 0.00 0.28 0.39
davinci 0.50 0.07 0.08 0.25 0.50 0.12 0.27 0.32
text-davinci-001 0.51 0.07 0.08 0.38 0.51 0.14 0.19 0.39
text-davinci-002 0.51 0.17 0.13 0.39 0.53 0.19 0.57 0.40
text-davinci-003 0.53 0.52 0.33 0.54 0.57 0.30 0.80 0.43
GPT-3.5-Turbo 0.51 0.38 0.40 0.44 0.58 0.30 0.73 0.51
GPT-4 0.55 0.60 0.31 0.74 0.67 0.42 0.91 0.46
GPT-4 (0-shot ICL) 0.60 0.19 0.25 0.72 0.65 0.27 0.85 0.48
GPT-4 (1-shot ICL) 0.66 0.24 0.30 0.70 0.71 0.38 0.78 0.41
GPT-4 (3-shot ICL) 0.61 0.70 0.70 0.75 0.69 0.29 0.70 0.40
GPT-4 (0-shot CoT) 0.57 0.57 0.28 0.73 0.66 0.43 0.90 0.53
GPT-4 (Manual CoT) 0.66 0.93 0.91 0.69 0.77 0.80 0.89 0.48
GPT-4 (EF) 0.60 - - 0.72 0.70 - 0.87 0.53
Table 3: Performance (accuracy) of different models in causal tasks in the ladder of causation (Rung 1 $\sim$ Rung 3) on different datasets, including CLADDER, CaLM, CRASS, and E-CARE. The column in gray correspond to tasks in Rung 1 (corr=correlation), the columns in white involve tasks in Rung 2 (ATE=average treatment effect; CDE = controlled direct effect; BAJ= backdoor adjustment); the columns in cyan correspond to tasks in Rung 3 (CR=counterfactual reasoning; NDE=natural direct effect; CE=causal explanation). In the upper part, we show results with the basic prompt; while in the lower part, we show results of GPT-4 with different prompting strategies. The results are collected from Chen et al. (2024a) and Jin et al. (2023a). Note that the experimental settings such as prompt templates may be different.
Causal discovery aims to identify the causal relationships between different variables, often serving as a fundamental step in real-world causal analysis. Most traditional causal discovery approaches rely on the data values and use statistical approaches to infer the underlying causal structure over the corresponding variables. These approaches include constraint-based methods (e.g., PC algorithm Spirtes et al. (2000) and FCI algorithm Spirtes et al. (2013); Zhang (2008)) which infer causal relationships by leveraging conditional independence tests, and score-based methods which assign scores to candidate causal graphs w.r.t. certain scoring criterion and seek the candidate causal graph with the highest score (e.g., GES algorithm Chickering (2002)). Various classical statistical approaches and recent machine learning or deep learning technologies Spirtes and Zhang (2016); Nogueira et al. (2022); Vowels et al. (2022) have been used in causal discovery.
Recent developments in LLMs provide new perspectives for causal discovery Takayama et al. (2024); Liu et al. (2024a); Kıcıman et al. (2023). Different from most existing causal discovery methods which can only utilize the data values of variables, LLMs can also leverage the metadata (e.g., the names of variables, the problem context) related to these variables to discover the implicit causal relationships. This reasoning process makes LLM-based causal discovery closer to human recognition. Recent literature Kıcıman et al. (2023) refer to this ability as knowledge-based causal discovery, and their experiments show that LLM-based knowledge-based causal discovery outperforms existing causal discovery methods on benchmarks Mooij et al. (2016). Currently, a variety of investigations have been conducted on LLMs in causal discovery tasks Kıcıman et al. (2023); Cai et al. (2023); Gao et al. (2023); Jin et al. (2023b); Long et al. (2023). These investigations are often conducted in the form of multi-choice or free-text question-answering, and they can mainly be divided into two types: pairwise causal discovery and full causal graph discovery.
Pairwise causal discovery (PCD) focuses on a pair of variables, either aiming to infer the causal direction ($A\rightarrow B$ or $A\leftarrow B$) between a given pair of variables $(A,B)$, or aiming to judge the existence of a causal relation between two variables. Among them, the experiments in Kıcıman et al. (2023) use the names of variables when constructing prompts, and their results show that LLMs (including GPT-3.5 and GPT-4) outperform state-of-the-art methods both on datasets with common variables (e.g., CauseEffectPairs Mooij et al. (2016)) and datasets that require particular domain knowledge (e.g., neuropathic pain Tu et al. (2019)). Despite the encouraging results, the empirical analysis from Zečević et al. (2023) implies that in many cases, LLMs are just “causal parrots" that repeat the embedded causal knowledge. A comparison between ChatGPT and fine-tuned small pre-trained language models Gao et al. (2023) shows LLMs’ advantage in some causal discovery tasks, but this work also discusses that the ability of LLMs in determining the existence of a causal relationship is worse than simply selecting the cause or effect of an input event from given options. Jin et al. (2023b) proposes a correlation-to-causation inference (Corr2Cause) task to evaluate the causal inference performance of LLMs. Their experimental results reveal that LLMs perform almost close to random on the task, even though this issue could be mitigated through fine-tuning, these models still have limitations in generalization on out-of-distribution settings.
Full causal graph discovery aims to identify the full causal graph that describes the causal relationships among a given set of variables. Compared with pairwise causal discovery, discovering the full causal graph is a more complicated problem as it involves more variables. In a preliminary exploration Long et al. (2023), GPT-3 shows good performance in discovering the causal graph with 3-4 nodes for well-known causal relationships in the medical domain. In more complicated scenarios, the ability of different versions of GPT to discover causal edges Kıcıman et al. (2023) has been validated on the neuropathic pain dataset Tu et al. (2019) with 100 pairs of true/false causal relations. LLM-based discovery (GPT-3.5 and GPT-4) on Arctic sea ice dataset Huang et al. (2021) has comparable or even better performance than representative baselines including NOTEARS Zheng et al. (2018) and DAG-GNN Yu et al. (2019). In Ban et al. (2023), the combination of the causal knowledge generated by LLMs and data-driven methods brings improvement in causal discovery in data from eight different domains with small causal graphs (5$\sim$48 variables and 4$\sim$84 causal relations). But similarly to PCD, LLMs also face many doubts and debates about their true ability in full graph discovery.
4.3 LLM for Causal Effect Estimation
Causal effect estimation aims to quantify how much manipulating a treatment can causally influence an outcome. In most cases, the causal effect of interest is estimated from observational data. Researchers in the NLP community have also made lots of efforts in causal effect estimation from text data Feder et al. (2022); Keith et al. (2020). Causal effect estimation on text data faces unique challenges due to the high-dimensional and complicated nature, for example, some important assumptions (e.g., positivity assumption D’Amour et al. (2021)) in traditional causal effect estimation are easily violated when high-dimensional text information is a confounder Keith et al. (2020). Fortunately, the NLP progress in recent decades, such as word embeddings Almeida and Xexéo (2019), topic modeling Blei et al. (2003) and dependency parsing Nivre (2005) have significantly contributed to estimating causal effects on text.
LLM has shown impressive performance in causal effect estimation as well. Recently, the connection between causal effect estimation and LLMs includes two different branches: (1) Causal effect in data: In this setting, LLMs aim to estimate the causal effect inside data Lin et al. (2023); Kıcıman et al. (2023) by leveraging their reasoning capability and properties (e.g., ability to handle large-scale training corpora). A benchmark for the capability of LLMs in causal inference, CLADDER Jin et al. (2023a), includes query types regarding causal effect estimation at different levels, e.g., ATE, ATT, NDE, NIE in the Rung 2, and ATT, NDE, NIE in the Rung 3. Existing evaluations show that the causal effect estimation task is still quite challenging for most LLMs. However, an encouraging finding is that proper techniques such as chain-of-thought (CoT) prompting strategy Jin et al. (2023a) can improve the performance significantly. (2) Causal effect in model: This setting aims to analyze the causal effect that involves the LLM itself. Most commonly, we focus on the causal effect of input data, model neurons, or learning strategies on LLMs’ predictions Vig et al. (2020); Meng et al. (2022); Stolfo et al. (2022). These studies can reveal the underlying LLM behavior and promote further investigations such as bias elimination Vig et al. (2020), model editing Meng et al. (2022), and robustness quantification Stolfo et al. (2022). For example, Stolfo et al. (2022) explores the causal effect of input (e.g., problem description and math operators) on output solutions in LLM-based mathematical reasoning. In Vig et al. (2020), a causal mediation analysis for gender bias propagated from model input to output is conducted in language models.
4.4 LLM for Other Causal Tasks
Experiments Chen et al. (2024a); Jin et al. (2023a); Kıcıman et al. (2023) have shown that there are various other causal inference tasks that LLMs can bring benefits to. (1) Causal attribution: LLMs show their capability in attribution tasks Kıcıman et al. (2023); Cai et al. (2023) typically in the forms of "why" or "what is the cause" questions. Related tasks also include identifying necessary or sufficient causes Liu et al. (2023); Kıcıman et al. (2023). By embedding human knowledge and cultural common sense, the results show that LLMs have the potential to flexibly address attribution problems in specific domains (such as law, economics, and medicine) where conventional methods may fall short Kıcıman et al. (2023). (2) Counterfactual reasoning: Recent studies Kıcıman et al. (2023); Jin et al. (2023a) conduct experiments on LLMs in different counterfactual reasoning scenarios, which are often in "what if" questions. While this task is one of the most challenging tasks in causal inference, the demonstrated improvement in LLMs compared to other methods is noteworthy. (3) Causal explanation: Many recent works investigate causal explanations based on queries on LLMs Bhattacharjee et al. (2023); Gat et al. (2023); Cai et al. (2023); Gao et al. (2023). Despite ongoing debates regarding LLM’s actual ability for causal reasoning, most empirical studies positively indicate that LLMs serve as effective causal explainers Gao et al. (2023). Such achievement is powered by LLMs’ capability of analyzing language logic and responding to questions using natural language.
4.5 Main Observations and Insights
From the evaluation discussed above and results shown in Table 2 and Table 3, we summarzie the main observations as follows: (1) Model performance: In general, many LLMs exhibit impressive performance in various causal tasks, especially in causal discovery, even with basic prompts. In some cases, their performance can be comparable to or even surpass human-level reasoning Kıcıman et al. (2023). However, as the task difficulty increases from Rung 1 to Rung 3, their performance becomes less satisfactory in higher-level complicated causal reasoning tasks Chen et al. (2024a). (2) Enhancement through proper strategies: The performance of LLMs can be significantly enhanced with effective prompting strategies, such as few-shot ICL and CoT. These approaches enable models to leverage contextual information and reasoning processes to achieve better results. Additionally, these models can provide valuable insights through causal explanations. (3) General patterns: While no definitive laws determine model performance universally, certain trends are still observable. For instance, scaling laws suggest that larger models generally perform better, although this is not always that straightforward. These trends provide valuable insights that can guide the future design and development of models. (4) Variability in model effectiveness: There is currently no universally superior LLM or strategy for causal tasks, as their effectiveness can vary significantly depending on the specific scenario. These observations highlight the need for more nuanced and adaptable approaches. (5) Common issues in LLMs for causal tasks: Current LLMs still struggle with many issues in causal tasks. For example, the answers often lack robustness and are highly sensitive to changes in prompts Kıcıman et al. (2023); Jin et al. (2023a). Besides, these models frequently default to memorizing and repeating information rather than engaging in actual causal reasoning Zečević et al. (2023), which can limit their effectiveness in complex causal scenarios. Besides, LLMs often fail to generate self-consistent answers for causal queries, i.e., the answers from LLMs often present causal relationships that conflict with each other. Ongoing debates and criticisms about whether LLM truly performs causal inference also compel more in-depth and precise analysis and evaluation.
5 Discussion and Future Prospects
In general, LLMs offer intriguing perspectives on causal inference, but current research also reveals many limitations, pointing to potential directions for future work that could advance the field Zhang et al. (2023); Kıcıman et al. (2023). Here, we outline several promising avenues for exploration in this research area, including:
  • Incorporating human knowledge: Many causal inference tasks necessitate human knowledge. A more comprehensive and intelligent integration of human knowledge (e.g., domain expertise) into LLMs could enhance causal reasoning, enabling interdisciplinary causal knowledge integration and analysis in both general settings and specialized domains such as finance, healthcare, and law Chen et al. (2024b).
  • Improving data generation: Real-world data often lack verified causal relations and counterfactuals. Utilizing LLMs for causal data generation can provide more diverse and realistic data, enriching real-world datasets with reliable causal relationships and improving training for causal reasoning models.
  • Addressing hallucinations: Hallucinations widely exist in LLM-generated causal inference answers. Focused efforts to eliminate hallucinations can lead to more accurate and reliable causal reasoning. Additionally, addressing broader issues such as fairness and biases can further enhance the trustworthiness of LLM in causal reasoning tasks.
  • Improving explanation and interactivity: Developing more interpretable and instructable LLMs in causal reasoning is crucial. Methods such as fine-tuning, model distillation, probing, and prompt engineering can improve how LLMs interact with users. Optimizing reasoning chains to accommodate human instructions and feedback can also foster more collaborative and controllable causal inference between humans and AI.
  • Exploring multimodal causality: Real-world scenarios often involve multiple modalities. Recent studies have begun exploring causality across different modalities, such as images Li et al. (2024) and videos Lam et al. (2024). Future research could further investigate these multimodal approaches to enhance causal understanding.
  • Developing a unified causal benchmark: There is currently no unified and widely recognized benchmark for evaluating causal performance in LLMs. Creating a comprehensive benchmark would facilitate better comparison and evaluation of different models, leading to more consistent and reliable assessments.
  • Addressing limitations in causal tasks and datasets: Existing causal tasks and benchmark datasets are relatively simplistic, often limited to multiple-choice and binary classification formats. These tasks frequently lack the complexity needed to address intricate scenarios and domain-specific contexts.
  • Advancing causality-focused models: Most existing methods only evaluate original LLMs without sufficient focus on causality-centric model designs. There is a significant opportunity for further research and development in this area to deepen models’ understanding of causality and improve their effectiveness in causal inference tasks.
6 Limitations
In this survey paper, it is important to acknowledge certain limitations that shape the scope and focus of our review. Firstly, our analysis is primarily centered on the application of large language models (LLMs) for causal inference tasks, thereby excluding exploration into how causality is utilized within LLM frameworks themselves. This decision provides a targeted perspective on leveraging LLMs to enhance causal inference methodologies but does not delve into the internal mechanisms or implementations of causal reasoning within these models.
Secondly, while we comprehensively examine the technical aspects and methodological advancements in using LLMs for causal inference, we do not extensively discuss ethical considerations or potential societal impacts associated with these applications. Ethical dimensions, such as fairness, bias mitigation, and privacy concerns, are critical in the deployment of AI technologies, including LLMs, and warrant dedicated attention and scrutiny in future research and applications. Addressing these limitations ensures a nuanced understanding of the opportunities and challenges in harnessing LLMs for causal inference while also advocating for responsible and ethical AI development and deployment practices.